OpenStudy (anonymous):

What is the sum of a 34–term arithmetic sequence where the first term is 14 and the last term is 146?

OpenStudy (anonymous):

do you know what is arithmetic sequence ?

OpenStudy (anonymous):

use the formula we gave in the last problem, tell us your answer and we'll check it :)

OpenStudy (anonymous):

to foolformath, not really. I am really confused with arithmetic sequence. I know one of the equations!

OpenStudy (anonymous):

The Sn=n/2(a1+an)

OpenStudy (anonymous):

http://www.purplemath.com/modules/series3.htm see if this helps

OpenStudy (anonymous):

Please don't memorize formulas .. understand it .. this is really elementary algebra ..

OpenStudy (anonymous):

Is it 2720?

OpenStudy (anonymous):

But I don't understand it..

OpenStudy (anonymous):

it can't be 2720 b/c the last term is 146.........

OpenStudy (anonymous):

S = 0.5n(a+l) S = 0.5(34)(14 +146) S = 0.5(34)(160) S = 17(160) S = 2720 oooopppppppsssssss u were right :)

OpenStudy (anonymous):

alright, thank you so much (: